OpenStudy (anonymous):

Need help transforming this polar coordinate: r^2=4cos(theta) into a parametric equation.

OpenStudy (anonymous):

\[r=\sqrt{x^2+y^2}\]

OpenStudy (anonymous):

\[x=rcos(\theta)\]

OpenStudy (anonymous):

multiply both sides by r that might elp

OpenStudy (anonymous):

\[r^3=4rcos(\theta)\] \[r^3=4x\]

OpenStudy (anonymous):

Oh I see you multiply each side by r...I guess I don't understand that part...

OpenStudy (anonymous):

it has to do with getting an r on the right side so you can simplify

OpenStudy (anonymous):

it or change it to cartesian

OpenStudy (anonymous):

o wait parametric hold on

OpenStudy (anonymous):

I just need to transform this equation so that I can graph it on a cartesian plane. I may have called it by the wrong name...

OpenStudy (anonymous):

yeah that's polar to cartesian so this is what ou need to know \[y=rsin(\theta)\] \[x=rcos(\theta)\] \[r=\sqrt{x^2+y^2}\]

OpenStudy (anonymous):

since you have \[r^2=4cos(\theta)\] you have no way of changing the right side because it has no r in the equation itself so by multiplying both sides youget \[r^3=4rcos(\theta)\] \[(\sqrt{x^2+y^2})^3=4x\]

OpenStudy (anonymous):

you write the left side as a exponential form \[(x^2+y^2)^{3/2}=4x\]

OpenStudy (anonymous):

now we want to get rid of the exponent by you can do that by multiply both sides by the reciprocal 3/2 \[x^2+y^2=(4x)^{3/2}\]

OpenStudy (anonymous):

Okay, cool thanks! I understand now :)

OpenStudy (turingtest):

nice! @Outkast3r09 but I have to point out the typo\[(x^2+y^2)^{3/2}=4x\implies x^2+y^2=(4x)^{2/3}\](exponent is reciprocal)

OpenStudy (anonymous):

And then graph, I solve for y, right. So, this would end up being... y=\[\sqrt{(4x) ^{2/3}-x ^{2}}\]

OpenStudy (turingtest):

I don't see why you need it solved for y, but if you did you would have to take \(\pm\) upon taking the sqrt you can also factor out an x^2\[y=\pm2x\sqrt{4x^{1/3}-1}\]but that seems unnecessary

OpenStudy (turingtest):

typo*\[y=\pm x\sqrt{(4x)^{1/3}-1}\]